Find the quartile deviation for the following set of 11 bottles of 1 litre of emery milk for


11 days by a dealer.


34, 37, 32, 25, 40, 45, 50, 42, 39, 47, 55


7. Calculate the quartile deviation and its coefficient from the following:


x 43 44 45 46 47 48 49


f 10 18 21 18 15 10 8

1
Expert's answer
2021-12-30T06:21:09-0500

6.


"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 x & f & cf \\ \hline
 25 & 1 & 1 \\
 \hdashline
 32 & 1 & 2 \\
 \hdashline
 34 & 1 & 3 \\
 \hdashline
 37 & 1 & 4 \\
 \hdashline
 39 & 1 & 5 \\
 \hdashline
 40 & 1 & 6 \\
 \hdashline
 42 & 1 & 7 \\
 \hdashline
 45 & 1 & 8 \\
 \hdashline
 47 & 1 & 9 \\
 \hdashline
 50 & 1 & 10 \\
 \hdashline
 55 & 1 & 11 \\
 \hdashline
\end{array}"

Here, "n=11."

"Q_1=(\dfrac{n+1}{4})" th value of the observation


"=(\dfrac{11+1}{4})" th value of the observation


"=(3)" th value of the observation


"=34"


"Q_3=(\dfrac{3(n+1)}{4})" th value of the observation


"=(\dfrac{3(11+1)}{4})" th value of the observation


"=(9)" th value of the observation


"=47"

Quartile deviation "=\dfrac{Q_3-Q_1}{2}=\dfrac{47-34}{2}=6.5"


7.


"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 x & f & cf \\ \hline
 43 & 10 & 10 \\
 \hdashline
 44 & 18 & 28 \\
 \hdashline
 45 & 21 & 49 \\
 \hdashline
 46 & 18 & 67 \\
 \hdashline
 47 & 15 & 82 \\
 \hdashline
 48 & 10 & 92 \\
 \hdashline
 49 & 8 & 100 \\
 \hdashline
\end{array}"

Here, "n=100."

"Q_1=(\dfrac{n+1}{4})" th value of the observation


"=(\dfrac{100+1}{4})" th value of the observation


"=(25.25)" th value of the observation


"=44"


"Q_3=(\dfrac{3(n+1)}{4})" th value of the observation


"=(\dfrac{3(100+1)}{4})" th value of the observation


"=(75.75)" th value of the observation


"=47"

Quartile deviation "=\dfrac{Q_3-Q_1}{2}=\dfrac{47-44}{2}=1.5"


Coefficient of Quartile deviation "=\dfrac{Q_3-Q_1}{Q_3+Q_1}=\dfrac{47-44}{47+44}=\dfrac{3}{91}\approx0.033"